# Query complexities of quantum channel discrimination and estimation: A unified approach

### v1

@@ 4.1.2 Bures distance of states
$d_{B}(\rho,\sigma)$ $\coloneqq\inf_{U\in\mathbb{U}}\left\|\sqrt{\rho}-U\sqrt{\sigma}\right\|_{2}.$ (4.23)
@@
$=\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sqrt{2\left(1-\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]\right)}$ (4.26)
@@ 4.2 Bures distance of channels
$=\sqrt{2}\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left\|I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right\|^{\frac{1}{2}},$ (4.41)
@@
$=\sup_{\left|\psi\right\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)},$ (4.49)
@@
$\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$
$=\sup_{\rho\in\mathbb{D},U\in\mathbb{U}}\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.56)
$=\sup_{\rho\in\mathbb{D}}\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.57)
$=\sqrt{2\left(1-\inf_{\rho\in\mathbb{D}}\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.58)
$=\sqrt{2\left(1-\sup_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\inf_{\rho\in\mathbb{D}}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.59)
$=\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sup_{\rho\in\mathbb{D}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.60)
$=\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$ (4.61)
@@ 6.1 Upper bound on Bures distance of channels in the parallel setting
$d_{B}^{2}\!\left(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}\right)\leq n\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left\|I-M_{W}\right\|^{2}\right\},$ (6.1)
@@ 6.2 Upper bound on Bures distance of channels in the adaptive setting
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right)\leq\\ n\inf_{W\in\mathbb{L}:\left\|W\right\|\leq 1}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M_{W}\right]\right\|\right)^{\frac{1}{2}}\left\|I-M_{W}\right\|\right\},$ (6.25)
@@
The upper bound in Theorem 16 does not exceed the upper bound in Theorem 18, consistent with the intuition that the Bures distance for parallel protocols should not exceed the Bures distance of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 16 and 18 and noting that the following inequality holds for every contraction $W$:
@@ 7.2 Upper bound on SLD Fisher information in the adaptive setting
The upper bound in Theorem 21 does not exceed the upper bound in Theorem 22, consistent with the intuition that the SLD Fisher information for parallel protocols should not exceed the SLD Fisher information of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 21 and 22 and noting that the following inequality holds for every Hermitian operator $H_{\theta}$ and isometry $V_{\theta}$:
@@ 8.1 Lower bounds on query complexities of channel discrimination
$1-\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\},$ (8.1)
$1-\frac{p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}\right\},$ (8.2)
@@
$n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil$ (8.15)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{L}}\left\{a_{W}:\left\|W\right\|\leq 1,\ M_{W}=I\right\}}\right\rceil,$ (8.16)
$n_{\operatorname{ad}}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\frac{c_{W}+\sqrt{c_{W}^{2}+4\sqrt{a_{W}b_{W}}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2\sqrt{a_{W}b_{W}}}\right\rceil$ (8.17)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{L}}\left\{a_{W}:\left\|W\right\|\leq 1,\ M_{W}=I\right\}}\right\rceil,$ (8.18)
@@
Applying Definition 1, let $n\in\mathbb{N}$ be such that $p_{e,\|}\leq\varepsilon$, where we have adopted the shorthand $p_{e,\|}$, as defined in (8.5). Let $W\in\mathbb{L}$ be a contraction (i.e., $\left\|W\right\|\leq 1$). Then from the monotonicity of $x\mapsto x\left(1-x\right)$ on $\left[0,\frac{1}{2}\right]$ and Corollary 24, we conclude that
@@
Since this lower bound holds for every contraction $W\in\mathbb{L}$, we can take a supremum over all such contractions to conclude that
$n\geq\sup_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\lceil\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil.$ (8.26)
@@
Since this lower bound holds for every contraction $W\in\mathbb{L}$ satisfying $M_{W}=I$, we can take a supremum over all such contractions (which is equivalent to an infimum over the denominator) to conclude that
$n\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{L}}\left\{a_{W}:\left\|W\right\|\leq 1,\ M_{W}=I\right\}}\right\rceil.$ (8.28)
@@
$\inf_{W\in\mathbb{L}}\left\{a_{W}:\left\|W\right\|\leq 1,\ M_{W}=I\right\}=+\infty,$ (8.29)
@@
$\inf_{W\in\mathbb{L}}\left\{a_{W}:\left\|W\right\|\leq 1,\ M_{W}=I\right\}=\\ \inf_{\lambda\geq 0,W\in\mathbb{L}}\left\{\lambda:\lambda I\geq 2\left(I-\Re\!\left[M_{W}\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M_{W}=I\right\},$ (8.30)
@@ 8.1.1 Finding lower bounds on query complexities of channel discrimination by binary search
$g(n)\coloneqq\inf_{\begin{subarray}{c}W\in\mathbb{L}:\\ \left\|W\right\|\leq 1\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\}.$ (8.38)
@@
Given this observation, we can find the smallest integer $n$ satisfying (8.37) (i.e., the query complexity $n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$) by binary search. Before doing so, it is helpful to have an upper bound on the query complexity. Toward this end, let us recall [NW25, Theorem 8], which implies the following upper bound on the query complexity $n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ by picking $s=\frac{1}{2}$ therein:
$n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\leq\left\lceil\frac{\ln\!\left(\frac{\sqrt{pq}}{\varepsilon}\right)}{Q_{\frac{1}{2}}(\mathcal{N}_{1}\|\mathcal{N}_{2})}\right\rceil,$ (8.40)
@@
The binary search algorithm for finding a lower bound on $n_{\|}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
@@
      Solve $g(n_{m})=\inf_{W\in\mathbb{L}}\left\{a_{W}+\left(n_{m}-1\right)b_{W}:\left\|W\right\|\leq 1\right\}$ by semi-definite optimization (see Proposition 38).
@@
The binary search algorithm for finding a lower bound on $n_{\operatorname{ad}}^{*}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
@@
      Solve $g_{\operatorname{ad}}(n_{m})=\inf_{W\in\mathbb{L}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}:\left\|W\right\|\leq 1\right\}$ by the procedure outlined in Remark 40.
@@
Let us finally note that both of these algorithms are applicable and useful to determine lower bounds on the minimum number of queries needed to achieve perfect channel discrimination (i.e., $\varepsilon=0$). Indeed, Ref. [DFY09] established conditions under which two channels can be discriminated perfectly. By setting $\varepsilon=0$ in Algorithms 28 and 29, we can thus determine lower bounds on the query complexities of perfect channel discrimination in both the parallel and adaptive access models, respectively.

### v2

@@ 4.1.2 Bures distance of states
$d_{B}(\rho,\sigma)$ $\coloneqq\inf_{U\in\mathbb{U}}\left\|\sqrt{\rho}-U\sqrt{\sigma}\right\|_{2}$ (4.23)
@@
$=\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi^{\rho}|\left(W\otimes I\right)|\psi^{\sigma}\rangle\right]\right)}$ (4.26)
@@ 4.2 Bures distance of channels
$=\sqrt{2}\inf_{W\in\mathbb{B}}\left\|I-\Re\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\right]\right\|^{\frac{1}{2}},$ (4.41)
@@
$=\sup_{\left|\psi\right\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)},$ (4.49)
@@
$\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$
$=\sup_{\rho\in\mathbb{D},U\in\mathbb{U}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.56)
$=\sup_{\rho\in\mathbb{D}}\inf_{W\in\mathbb{B}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.57)
$=\sqrt{2\left(1-\inf_{\rho\in\mathbb{D}}\sup_{W\in\mathbb{B}}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.58)
$=\sqrt{2\left(1-\sup_{W\in\mathbb{B}}\inf_{\rho\in\mathbb{D}}\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.59)
$=\inf_{W\in\mathbb{B}}\sup_{\rho\in\mathbb{D}}\sqrt{2\left(1-\Re\!\left[\operatorname{Tr}\!\left[V_{1}^{{\dagger}}\left(W_{E}\otimes I_{B}\right)V_{2}\rho\right]\right]\right)}$ (4.60)
$=\inf_{W\in\mathbb{B}}\sup_{|\psi\rangle_{RA}\in\mathbb{P}}\sqrt{2\left(1-\Re\!\left[\langle\psi|_{RA}V_{1}^{{\dagger}}\left(W_{E}\otimes I_{RB}\right)V_{2}|\psi\rangle_{RA}\right]\right)}$ (4.61)
@@ 6.1 Upper bound on Bures distance of channels in the parallel setting
$d_{B}^{2}\!\left(\mathcal{N}_{1}^{\otimes n},\mathcal{N}_{2}^{\otimes n}\right)\leq n\inf_{W\in\mathbb{B}}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left\|I-M_{W}\right\|^{2}\right\},$ (6.1)
@@ 6.2 Upper bound on Bures distance of channels in the adaptive setting
$\sup_{\left(\mathcal{A}_{i}\right)_{i=1}^{n-1}}d_{B}^{2}\!\left(\mathcal{P}_{1}^{(n)},\mathcal{P}_{2}^{(n)}\right)\leq\\ n\inf_{W\in\mathbb{B}}\left\{2\left\|I-\Re\!\left[M_{W}\right]\right\|+\left(n-1\right)\left(2\left\|I-\Re\!\left[M_{W}\right]\right\|\right)^{\frac{1}{2}}\left\|I-M_{W}\right\|\right\},$ (6.25)
@@
The upper bound in Theorem 17 does not exceed the upper bound in Theorem 19, consistent with the intuition that the Bures distance for parallel protocols should not exceed the Bures distance of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 17 and 19 and noting that the following inequality holds for every contraction $W$:
@@ 7.2 Upper bound on SLD Fisher information in the adaptive setting
The upper bound in Theorem 22 does not exceed the upper bound in Theorem 23, consistent with the intuition that the SLD Fisher information for parallel protocols should not exceed the SLD Fisher information of adaptive protocols, given that parallel protocols are contained with the set of all adaptive protocols. This can be seen by comparing the objective functions in Theorems 22 and 23 and noting that the following inequality holds for every Hermitian operator $H_{\theta}$ and isometry $V_{\theta}$:
@@ 8.1 Lower bounds on query complexities of channel discrimination
$1-\frac{p_{e,\|}\left(1-p_{e,\|}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\},$ (8.1)
$1-\frac{p_{e,\operatorname{ad}}\left(1-p_{e,\operatorname{ad}}\right)}{pq}$ $\leq n\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}\right\},$ (8.2)
@@
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{W\in\mathbb{B}}\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil$ (8.15)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil,$ (8.16)
$n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\geq\left\lceil\sup_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\frac{c_{W}+\sqrt{c_{W}^{2}+4\sqrt{a_{W}b_{W}}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2\sqrt{a_{W}b_{W}}}\right\rceil$ (8.17)
$\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil,$ (8.18)
@@
Applying Definition 1, let $n\in\mathbb{N}$ be such that $p_{e,\|}\leq\varepsilon$, where we have adopted the shorthand $p_{e,\|}$, as defined in (8.5). Let $W\in\mathbb{B}$ be a contraction (i.e., $\left\|W\right\|\leq 1$). Then from the monotonicity of $x\mapsto x\left(1-x\right)$ on $\left[0,\frac{1}{2}\right]$ and Corollary 25, we conclude that
@@
Since this lower bound holds for every contraction $W\in\mathbb{B}$, we can take a supremum over all such contractions to conclude that
$n\geq\sup_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\lceil\frac{b_{W}-a_{W}+\sqrt{\left(b_{W}-a_{W}\right)^{2}+4b_{W}\left(1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}\right)}}{2b_{W}}\right\rceil.$ (8.26)
@@
Since this lower bound holds for every contraction $W\in\mathbb{B}$ satisfying $M_{W}=I$, we can take a supremum over all such contractions (which is equivalent to an infimum over the denominator) to conclude that
$n\geq\left\lceil\frac{1-\frac{\varepsilon\left(1-\varepsilon\right)}{pq}}{\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}}\right\rceil.$ (8.28)
@@
$\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}=+\infty,$ (8.29)
@@
$\inf_{W\in\mathbb{B}}\left\{a_{W}:M_{W}=I\right\}=\\ \inf_{\lambda\geq 0,W\in\mathbb{L}}\left\{\lambda:\lambda I\geq 2\left(I-\Re\!\left[M_{W}\right]\right)\geq-\lambda I,\ \begin{bmatrix}I&W\\ W^{{\dagger}}&I\end{bmatrix}\geq 0,\ M_{W}=I\right\},$ (8.30)
@@ 8.1.1 Finding lower bounds on query complexities of channel discrimination by binary search
$g(n)\coloneqq\inf_{\begin{subarray}{c}W\in\mathbb{B}\end{subarray}}\left\{a_{W}+\left(n-1\right)b_{W}\right\}.$ (8.38)
@@
Given this observation, we can find the smallest integer $n$ satisfying (8.37) (i.e., the query complexity $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$) by binary search. Before doing so, it is helpful to have an upper bound on the query complexity. Toward this end, let us recall [NW25, Theorem 8], which implies the following upper bound on the query complexity $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ by picking $s=\frac{1}{2}$ therein:
$n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ $\leq\left\lceil\frac{\ln\!\left(\frac{\sqrt{pq}}{\varepsilon}\right)}{Q_{\frac{1}{2}}(\mathcal{N}_{1}\|\mathcal{N}_{2})}\right\rceil,$ (8.40)
@@
The binary search algorithm for finding a lower bound on $n_{\|}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
@@
      Solve $g(n_{m})=\inf_{W\in\mathbb{B}}\left\{a_{W}+\left(n_{m}-1\right)b_{W}\right\}$ by semi-definite optimization (see Proposition 39).
@@
The binary search algorithm for finding a lower bound on $n_{\operatorname{ad}}^{\star}(p,\mathcal{N}_{1},q,\mathcal{N}_{2},\varepsilon)$ proceeds according to the following steps:
@@
      Solve $g_{\operatorname{ad}}(n_{m})=\inf_{W\in\mathbb{B}}\left\{a_{W}+\left(n-1\right)\sqrt{a_{W}b_{W}}\right\}$ by the procedure outlined in Remark 41.
@@
Let us finally note that both of these algorithms are applicable and useful to determine lower bounds on the minimum number of queries needed to achieve perfect channel discrimination (i.e., $\varepsilon=0$). Indeed, Ref. [DFY09] established conditions under which two channels can be discriminated perfectly. By setting $\varepsilon=0$ in Algorithms 29 and 30, we can thus determine lower bounds on the query complexities of perfect channel discrimination in both the parallel and adaptive access models, respectively.